OpenStudy (anonymous):

The product of a number and -7/10 is 1/2 . What is the number? -5/7 5/7 -7/20 7/20

OpenStudy (anonymous):

State it as an algebra problem:

OpenStudy (anonymous):

x * -7/10 = 1/2

OpenStudy (anonymous):

Multiply both sides by 10/-7

OpenStudy (anonymous):

does it help if we write your problem in this form \[-\frac{7}{10}*x=\frac{1}{2}\]

OpenStudy (anonymous):

so then we get x * -7/10 * 10/-7 = 1/2 * 10/-7 and [-7/10 * 10/-7] = 1 so we get x = 1/2 * 10/-7

OpenStudy (anonymous):

and we multiply numerators and denominators together when multiplying fractions, so you get 10/-14 which we can reduce to 5/-7 and we can move the negative to the numerator if we want x= -5 / 7 Which is about equal to -.714

OpenStudy (anonymous):

so its -5/7?

OpenStudy (anonymous):

no 5/7 not -5/7

OpenStudy (anonymous):

It is -5/7

OpenStudy (anonymous):

negative times a positive = negative negative times a negative = positive

OpenStudy (anonymous):

Two negative numbers multiply to make a positive

OpenStudy (anonymous):

thanks :)